A 30.0 mL sample of 0.150 M H2so4 is placed into 12 mL Erlenmeyer flask and a 25 mL buret is filled with 0.250 M KoH. Calculate the volume of base required to reach the equivalence point you must not use M2V2 calculation.
Answer #1

no. of mole = molarity \times volume of solution in liter

no. of mole of H2SO4 = 0.150 \times 0.030 = 0.0045 mole

neutrilization reaction between KOH and H2SO4 is

2KOH + H2SO4  \rightarrow  2H2O + K2SO4

According to reaction 2 mole of KOH neutilize 1 mole of H2SO4 therefore to neutrilize 0.0045 mole of H2SO4 require KOH = 0.0045 \times 2 = 0.0090 mole of KOH

volume of solution in liter = no.of mole / molarity

volume of NaOH = 0.0090 / 0.250 = 0.036 liter = 36 ml

36 ml NaOH required